OpenStudy (anonymous):

write the equation 2x^2+2y^2=8x-16y-32

OpenStudy (anonymous):

state whether the graph is a parabola, ellipse,hyperbola or circle

OpenStudy (anonymous):

It is a circle.

OpenStudy (campbell_st):

2x^2 - 8x + 2y^2 +16y = -32 x^2 - 4x + y^2 + 8y = -16 x^2 - 4x +y^2 + 8y + 16 = 0 not a lot else to do...

OpenStudy (anonymous):

2 (x-2)^2+2 (y+4)^2-8 = 0

OpenStudy (dumbcow):

Divide out a 2 --> (x-2)^2 +(y+4)^2 = 4 circle with radius 2 and center (2,-4)